1. Calculate the pH of a 1.08×10-2 M solution of the decongestant ephedrine hydrochloride if the pKb of ephedrine (its conjugate base) is 3.86.
1
Expert's answer
2020-04-14T01:43:17-0400

The acid-base equilibrium is:

EphH+ = Eph + H+

The constant is Ka = Kw/Kb, where pKw = 14,

pKa = 14 - pKb = 10.14.

The equilibrium concentration of [H+] is equal to [Eph]. The equilibrium concentration of [EphH+] = 1.08*10-2 - [H+].

The equilibrium constant is very small (10-10.14) and the concentration of EphH+ is by 8 orders of magnitude higher (1.08*10-2), that is why the degree of dissociation is negligible and the equilibrium concentration of [EphH+] is close its initial concentration 1.08*10-2.

Thus, Ka = 10-10.14 = [H+]2/[EpH+] = [H+]2/(1.08*10-2),

in terms of logarithms:

pH = (pKa + 2 - lg1.08)/2 = 6.05